The difference between the area of a regular hexagon of side 72cm and area of its inscribed circle is?

Grade:10

1 Answers

Vikas TU
4 years ago
Diameter of the inscribed circle in a regular polygon of n number of side (6, in this case) each of length, s (72 m in this case) = Dic = s*cot(180/n) = 72 *cot(180/6) = 72*cot 30 = 124.7050808 cm
Area of the inscribed circle = Aic = [(22/7)* 124.7050808^2]/4= 97.7142857 sq cm.
Diameter of the circumscribed circle in a regular polygon of n number of side (6, in this case) each of length, s (72 cm in this case) = Doc = s*cosec(180/n) = 72 *cosec(180/6) = 72*cosec 30 = 144 cm
Area of the circumscribed circle = Aoc = [(22/7)* 144^2]/4 = 16286.2857142 sq cm.